OpenStudy (anonymous):

check my answer what is the slope of the line passing through the point (4 -2) and perpendicular to the line y=4x+2?write your answer as a fraction.

OpenStudy (anonymous):

i think its -1/4

OpenStudy (anonymous):

im not so sure but i think thats right

OpenStudy (mendicant_bias):

Yeah, I think you're right, I'm wondering why they included two pieces of information when only one was necessary. From the point-slope equation, you can find the slope, which you obviously have, and that should be it. It's not asking for an intercept, so I don't know why it gave you that point. You can find the slope of the new equation by just plugging in the x and y values of the point that your new line passes through and making them equal the slope, e.g. \[y = -\frac{ 1 }{ 4 }x + b\] \[-2 = -\frac{ 1 }{ 4 }(4) + b\] \[-1 = b\] So I think you're just plain right, but the question itself was poorly written.